Six utility men can clean the entire three star hotel for 10 hours, If the hotel needs to be clean for 6 hours, how many utility men are needed to finish the work


1
Expert's answer
2021-11-03T15:13:53-0400

The speed of 1 worker is


"\dfrac{1}{6(10)} hotel/hour"

The number of workers is


"n=\dfrac{1\ hotel}{6\ hour(\dfrac{1}{6(10)} hotel/hour)}=10"

10 men are needed  to finish the work.